# AgriGPT-VL: Agricultural Vision–Language Understanding Suite

### v2

@@ 4.1 Comparative Experiment
Figure 8: Multimodal evaluation of vision-language models on AgriBench-VL-4K.
@@
Table 2: Vision–language evaluation (multimodal capability). Comparison of AgriGPT-VL with general VLMs on image-grounded tasks.
@@
We focus on two questions: (i) after progressively injecting domain knowledge and vision–language alignment, is textual competence preserved and strengthened; and (ii) in real image–language settings, does the model exhibit stronger visual grounding and agronomic reasoning—i.e., can it both choose correctly (discriminative robustness) and articulate evidence–based answers (generation quality). To this end, we evaluate text–only capability on AgriBench-13K (Yang & et al., 2025)and multimodal capability on AgriBench-VL-4K. For discriminative evaluation, we report Acc (single–choice accuracy, scored per question) and Acc+ (image–level cross–consistency: both single–choice questions for the same image must be correct). For generation, we report BLEU (Papineni et al., 2002), METEOR (Banerjee & Lavie, 2005), and ROUGE-L (Lin, 2004) to measure terminology conformity, semantic coverage, and structural completeness.
@@
We compare AgriGPT-VL against twelve representative vision–language models: InternVL-3-8B/3-14B (Zhu et al., 2025), LLaVA-1.5-7B (Liu et al., 2023a), MiniCPM-V-2.6 (Yao et al., 2024; OpenBMB Team, 2024), Yi-VL-6B and Yi-VL-34B (01.AI, 2024b; a), Qwen-VL-7B (Bai et al., 2023), Qwen2.5-VL-72B-Instruct (Bai et al., 2025), Qwen-QVQ (Qwen Team, 2024), DeepSeek-VL-1.2 (Lu et al., 2024), and Gemini-2.5 Flash/Pro (Google DeepMind & Google AI, 2025a; b).
@@
As shown in tables 1 and figure 6, on AgriBench-13K (Yang & et al., 2025), AgriGPT-VL leads across mainstream text metrics, indicating that the progressive training does not sacrifice language ability; instead, it strengthens standardized use of agricultural terminology and canonical answer style, consolidating textual representations and providing a stable linguistic base for the subsequent multimodal stage.
@@
As shown in table 2 and figure 8, on AgriBench-VL-4K, we obtain the best results on all metrics, surpassing several flagship large models. Gains in Acc reflect more precise image–option matching; gains in Acc+ demonstrate consistent semantics per image and stronger resistance to hard distractors (confusable taxa and co–occurring conditions), thereby mitigating chance guessing and better reflecting true capability. Improvements in Bleu (Papineni et al., 2002), Meteor (Banerjee & Lavie, 2005), and Rouge-1-f/Rouge-2-f/Rouge-L-f (Lin, 2004) further indicate three strengthened abilities: (1) visual evidence grounding and factor extraction (organs, colors/lesions, phenology); (2) agronomic multi-step reasoning (from symptoms to plausible causes and management consistent with scene constraints); and (3) professional, audit-ready expression (units, terminology, and thresholds that follow domain conventions).Detailed definitions and computation formulas of the evaluation metrics are included in Appendix A.5
@@
In addition, as shown in figure 7, we conduct JudgeLM (Jiao et al., 2023) blind pairwise comparisons: for each query, two systems’ outputs are judged head-to-head, we swap left/right positions to reduce order bias, and average the two outcomes. We report three preference metrics: WR (ties excluded). Across most strong baselines, AgriGPT-VL achieves consistently higher win rates and remains competitive against top large models, corroborating the above advantages from a preference perspective. Appendix A.3 describes the prompt design for the LLM-based judger, and Appendix A.4 details the metric computation methodology.
@@ 4.2 Ablation Study
Figure 9: Ablation study on curriculum-based alignment stages of AgriGPT-VL
Table 3: Ablation study of alignment stages. Bold indicate best per column.
@@
Table 4: Evaluation of general capabilities before and after fine-tuning. Numbers in the second row indicate dataset sample sizes.
@@
As shown in tables 3 and figure 9, the results reveal a clear hierarchy of contributions: Shallow Alignment primarily improves lexical and descriptive consistency, stabilizing image–text keypoint correspondence; Deep Alignment is the main driver of cross-modal understanding and reasoning, lifting both discriminative and generation metrics; and GRPO further enhances factual faithfulness and robustness, with the largest gains on the stricter image-level cross-consistency metric (Acc+), indicating that expert-level instructions are necessary to constrain high-precision behavior.
@@ 4.3 Generalization Evaluation
To assess whether domain specialization preserves general capabilities, we compare the fine-tuned AgriGPT-VL with its base model (Qwen2.5-VL) on six public benchmarks: three text-only (MMLU (Hendrycks et al., 2021), ARC (Clark et al., 2018), OpenBookQA (Mihaylov et al., 2018)) and three vision–language (MMBench (Liu et al., 2025), MMMU (Yue et al., 2024), SEED-Bench (Li et al., 2024)).

### v3

@@ 4.1 Comparative Experiment
Table 2: Performance comparison on the AgriBench-VL-4K benchmark. Acc and Acc+ correspond to accuracy on single-choice visual reasoning questions, whereas the other metrics evaluate open-ended question–answering quality.
@@
We focus on two questions: (i) after progressively injecting domain knowledge and vision–language alignment, is textual competence preserved and strengthened; and (ii) in real image–language settings, does the model exhibit stronger visual grounding and agronomic reasoning—i.e., can it both choose correctly (discriminative robustness) and articulate evidence–based answers (generation quality). To this end, we evaluate text–only capability on AgriBench-13K [Yang2025AgriGPT]and multimodal capability on AgriBench-VL-4K. For discriminative evaluation, we report Acc (single–choice accuracy, scored per question) and Acc+ (image–level cross–consistency: both single–choice questions for the same image must be correct). For generation, we report BLEU [papineni2002bleu], METEOR [banerjee2005meteor], and ROUGE-L [lin2004rouge] to measure terminology conformity, semantic coverage, and structural completeness.
@@
We compare AgriGPT-VL against twelve representative vision–language models: InternVL-3-8B/3-14B [zhu2025internvl3], LLaVA-1.5-7B [liu2023llava15], MiniCPM-V-2.6 [yao2024minicpmv, openbmb2024minicpmv26], Yi-VL-6B and Yi-VL-34B [yi2024yivl6b, yi2024yivl34b], Qwen-VL-7B [bai2023qwenvl], Qwen2.5-VL-72B-Instruct [bai2025qwen25vl], Qwen-QVQ [qwen2024qvq], DeepSeek-VL-1.2 [lu2024deepseekvl], and Gemini-2.5 Flash/Pro [google2025gemini25flash, google2025gemini25pro].
@@
As shown in tables 1, on AgriBench-13K [Yang2025AgriGPT], AgriGPT-VL leads across mainstream text metrics, indicating that the progressive training does not sacrifice language ability; instead, it strengthens standardized use of agricultural terminology and canonical answer style, consolidating textual representations and providing a stable linguistic base for the subsequent multimodal stage.
@@
As shown in table 2, on AgriBench-VL-4K, we obtain the best results on all metrics, surpassing several flagship large models. Gains in Acc reflect more precise image–option matching; gains in Acc+ demonstrate consistent semantics per image and stronger resistance to hard distractors (confusable taxa and co–occurring conditions), thereby mitigating chance guessing and better reflecting true capability. Improvements in Bleu [papineni2002bleu], Meteor [banerjee2005meteor], and Rouge-1-f/Rouge-2-f/Rouge-L-f [lin2004rouge] further indicate three strengthened abilities: (1) visual evidence grounding and factor extraction (organs, colors/lesions, phenology); (2) agronomic multi-step reasoning (from symptoms to plausible causes and management consistent with scene constraints); and (3) professional, audit-ready expression (units, terminology, and thresholds that follow domain conventions).Detailed definitions and computation formulas of the evaluation metrics are included in Appendix A.5
@@
In addition, as shown in figure 6, we conduct JudgeLM [jiao2023panda] blind pairwise comparisons: for each query, two systems’ outputs are judged head-to-head, we swap left/right positions to reduce order bias, and average the two outcomes. We report three preference metrics: WR (ties excluded). Across most strong baselines, AgriGPT-VL achieves consistently higher win rates and remains competitive against top large models, corroborating the above advantages from a preference perspective. Appendix A.3 describes the prompt design for the LLM-based judger, and Appendix A.4 details the metric computation methodology.
@@ 4.2 Ablation Study
Table 3: Ablation study of alignment stages. Bold indicates best per column.
@@
Table 4: Evaluation of general capabilities before and after fine-tuning.
@@
As shown in tables 3, the results reveal a clear hierarchy of contributions: Shallow Alignment primarily improves lexical and descriptive consistency, stabilizing image–text keypoint correspondence; Deep Alignment is the main driver of cross-modal understanding and reasoning, lifting both discriminative and generation metrics; and GRPO further enhances factual faithfulness and robustness, with the largest gains on the stricter image-level cross-consistency metric (Acc+), indicating that expert-level instructions are necessary to constrain high-precision behavior.
@@ 4.3 Generalization Evaluation
To assess whether domain specialization preserves general capabilities, we compare the fine-tuned AgriGPT-VL with its base model (Qwen2.5-VL) on six public benchmarks: three text-only (MMLU [hendrycks2021mmlu], ARC [clark2018arc], OpenBookQA [mihaylov2018can]) and three vision–language (MMBench [liu2025mmbench], MMMU [Yue_2024_CVPR], SEED-Bench [Li_2024_CVPR]).
@@
4.4 External Evaluation on AgriBench
To further assess the robustness of our model, we conduct an additional evaluation on AgriBench[zhou2024agribench], an external benchmark comprising approximately 700 single-choice questions. Unlike our in-house AgriBench-VL-4K, AgriBench differs entirely in data sources, annotation protocols, textual style, and question distribution, and is therefore fully decoupled from our training data and data-generation pipeline. This provides a genuinely out-of-distribution setting for examining whether the model can transfer across construction paradigms. Such an external evaluation not only eliminates potential biases arising from stylistic alignment but also offers a more comprehensive measure of the model’s ability to interpret real agricultural scenarios.
As shown in tables 5, AgriGPT-VL achieves the highest accuracy of (70.10%) on AgriBench[zhou2024agribench], clearly outperforming publicly available vision–language models including Gemini-2.5-Pro (56.94%). The significance of this independent evaluation lies in demonstrating that the improvements of AgriGPT-VL do not rely on the stylistic patterns of our training data nor on overfitting to a single construction pipeline. Instead, the model exhibits stronger grounding in agricultural knowledge, pest and disease characteristics, and field-level visual cues. For agriculture, achieving clear gains on a fully external benchmark indicates robust cross-source generalization.
Table 5: Model accuracy on AgriBench dataset.
Model Acc
InternVL-3-8B 65.66%
InternVL-3-14B 60.08%
LLaVA-1.5-7B 59.29%
Qwen2.5-VL-7B 49.94%
Qwen2.5-VL-72B-Instruct 55.64%
Qwen/QVQ-72B-Preview 45.39%
DeepSeek-VL2 50.84%
Gemini-2.5-Flash 53.70%
Gemini-2.5-Pro 56.94%
AgriGPT-VL (ours) 70.10%